3) Assume that IQ scores of college graduates are normally distributed. A researcher collects a random...

3)

Assume that IQ scores of college graduates are normally distributed. A researcher collects a random sample of 24 college graduates and tests their IQ. The sample had a mean IQ of 109 with a standard deviation of 15.9.

a) Find a 95% confidence interval for the true mean IQ for college graduates.

b) Provide the right endpoint of the interval as your answer.

Round your answer to 1 decimal place.
Answer #1

a)

sample mean 'x̄= 109.000
sample size   n= 24.00
sample std deviation s= 15.900
std error 'sx=s/√n= 3.2456
for 95% CI; and 23 df, value of t= 2.069
margin of error E=t*std error    = 6.715
lower bound=sample mean-E = 102.28
Upper bound=sample mean+E = 115.72

b)

right endpoint of the interval =115.7